Jamie has 105 pieces of candy leftover from Halloween. She would like to distribute them evenly to the 7 kids on her block. Write an equation to show how many pieces of candy each kid will receive.


1
Expert's answer
2022-01-24T15:52:15-0500

Solution.

Jamie will evenly divide the candy left in her after Halloween, if you use the equation

"x=\frac{m}{n},"


where "x" is the number of candies that each child will receive, "m" is the number of candies that Jamie has, "n" is the number of children.

We will have


"x=\frac{105}{7}=15."

So, each child will receive 15 candies.

Answer. "x=\frac{105}{7}, x=15."